OpenStudy (anonymous):

help im confuse simplify each sum or difference 9. sqrt 5 + 6 sqrt 5 11. 7 sqrt 3 + sqrt 3 13. 3 sqrt 7 - sqrt 63

OpenStudy (anonymous):

For numbers 9 and 11, you can factor out the square root, ie \[9) \ \ 1\sqrt{5}+6\sqrt{5} = (1+6)\sqrt{5}\]

OpenStudy (anonymous):

For the last one, you can simplify square root 63 \[\sqrt{63}=\sqrt{7*6}=\sqrt{7}\sqrt{6}\] so your final "simplification" will have the square root 7 factored out like up above ^_^ Follow?

OpenStudy (anonymous):

im following

OpenStudy (anonymous):

so! what would the answer to number 9 be? ^^

OpenStudy (anonymous):

5sqrt5

OpenStudy (anonymous):

?

OpenStudy (anonymous):

no 7 sqrt 5

OpenStudy (anonymous):

yup yup !! Then how about #11?

OpenStudy (anonymous):

im not sure

OpenStudy (anonymous):

what could you factor out of those two terms like we did for #9?

OpenStudy (shamil98):

63 = 9 x 7 not 7 x 6

OpenStudy (anonymous):

i only got 42 so far

OpenStudy (shamil98):

\[\sqrt{63}=\sqrt{9} \times \sqrt{7}\]

OpenStudy (anonymous):

@shamil98 I REALLY should learn how to multiply one of these days :P Thanks!!

OpenStudy (anonymous):

@blubbernuggin How did you get 42?

OpenStudy (anonymous):

oh, 6*7. Yeah, I'm sorry, that was my bad. Brain hiccup.

OpenStudy (anonymous):

But that was for #13. Did you get 11 already?

OpenStudy (anonymous):

im working on it. i think

OpenStudy (anonymous):

hold on i think i got it

OpenStudy (anonymous):

Also, I really do apologize for factoring that wrong :/

OpenStudy (anonymous):

its ok i can multiply either so here is what i did \[3\sqrt{7} - \sqrt{63} = 3\sqrt{7} - \sqrt{9 x 7} = 3\sqrt{7} - \sqrt{9} x \sqrt{7} = (3-3) \sqrt{7} =\sqrt{7}\]

OpenStudy (anonymous):

cant*

OpenStudy (anonymous):

number 13

OpenStudy (anonymous):

close - you got the hard part. What's \[3-3=?\] ?

OpenStudy (anonymous):

3 - 3 is 0

OpenStudy (anonymous):

yah, so \[(3-3)\sqrt{7} = 0\sqrt{7} = 0\] ^_^

OpenStudy (anonymous):

oooooooooooohhhhhhh. <.<

OpenStudy (anonymous):

thank you i get it now

OpenStudy (anonymous):

very welcome ^_^

OpenStudy (anonymous):

hey wait can you help me one more time

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

\[3\sqrt{45} - 8\sqrt20 = (3-8) \]

OpenStudy (anonymous):

that all i got up to i dunno what to do after that

OpenStudy (anonymous):

unfortunately you can't factor out the square root if it's not the *same square root, so you can't pull the 3 and the -8 out just yet. You have to find a common factor in the radicals What's are common factors of 20 and 45?

OpenStudy (anonymous):

1 and 5 ?

OpenStudy (anonymous):

noo just 5

OpenStudy (anonymous):

yup yup, so what do you get when you factor out 5 from both of those? (like factoring the 63 into 6 and 9 in #13)

OpenStudy (anonymous):

you mean 9 and 7 ? lol

OpenStudy (anonymous):

0_o I do, I promise...

OpenStudy (anonymous):

but i think 25? or is it 9 and 4?

OpenStudy (anonymous):

Just checking to make sure you're paying attention, right? :/ :P

OpenStudy (anonymous):

9 and 4 are correct! So then you can finally factor out that square root 5 to get \[3\sqrt{9}\sqrt{5}-8\sqrt{4}\sqrt{5}=(3\sqrt{9}-8\sqrt{4})\sqrt{5}\] which you can simplify a little bit further

OpenStudy (anonymous):

so is the next step 27 - 32?

OpenStudy (anonymous):

those are still square roots of those numbers, so it'd be \[3*3-8*2\]

OpenStudy (anonymous):

so 9 - 16?

OpenStudy (anonymous):

= -7\[= -7\sqrt{5}\]

OpenStudy (anonymous):

yup yup yup ^_^

OpenStudy (anonymous):

or is it -35

OpenStudy (anonymous):

oh ok xD

OpenStudy (anonymous):

we've gotten all conflabulated! :P -7 is right though ^_^

OpenStudy (anonymous):

do the same steps apply with addition like for this problem?

OpenStudy (anonymous):

like if it were 3+8 instead of 3-8?

OpenStudy (anonymous):

if it were like this \[3\sqrt{45 }+ 8\sqrt{20}\]

OpenStudy (anonymous):

do same steps still apply even though the symbol changed

OpenStudy (anonymous):

most definitely. Factoring out the square root 5 has no effect on that sign, so the steps would be exactly the same.

OpenStudy (anonymous):

alright cool i got this thanks now beyond you go helping more kids indeed of help!

OpenStudy (anonymous):

in need*

OpenStudy (anonymous):

^_^ nice workin' with ya, and holler if you need anything else!